OpenStudy (anonymous):

if there are 8.800 x 10^25 atoms of tin, how many moles of tin is there and what would the mass be?

sam (.sam.):

\[8.800 \times 10^{25}~ atoms~Sn \times \frac{1 moles~Sn}{6.02 \times 10^{23}~ atoms~Sn}=146.2~moles~Sn\] \[146.2~moles~Sn \times \frac{118.7gSn}{1~moles~Sn}=17354g~Sn\]

OpenStudy (anonymous):

thank you so much! aha i owe you! <3

sam (.sam.):

hope you know how to do it ;)